Feed aggregator

iocipm5c00 restored

Recent Log Entries -

Log Number: 4153488Logbook(s): Tags: Body: restorelib 2.6 Target restore date was 2023-03-24 10:00:23 The output from the myRestore execution was as follows: iocipm5c00 <NOTE> bpmdr: Provided macro *IOC* is never used bpmdr bpmdr_ALPHA@2023-03-24 10:00 2(2) restored bpmdr_CALCONS@2023-03-24 10:00 17(17) restored bpmdr_CALFACS1@2023-03-24 10:00 6(6) restored bpmdr_CALFACS2@2023-03-24 10:00 12(12) restored bpmdr_DISBPM@2023-03-24 10:00 1(1) restored bpmdr_GOF2BPM@2023-03-24 10:00 12(12) restored bpmdr_GOFBPM@2023-03-24 10:00 5(5) restore

COMMENT ON Inj. 1/4 CM - Booster Swap

Recent Log Entries -

Log Number: 4153410Logbook(s): Body: 

Last 2 remaining girders removed, the 2 girder stands and their base plates have been removed, the girder pedestal was also removed yesterday, and all adjuster cartridges have been removed from the girder stands and pedestal. Area has been swept and there are no obstructions on the floor.

COMMENT ON Modify 0L02 RF drive system for booster upgrade

Recent Log Entries -

Log Number: 4153403Logbook(s): Body: 

Larry Farrish "This morning replaced all waveguide air supply hoses in the new zones NL21-26, except the main supply at the top of the ceiling I will get a taller ladder for that next week.
Waveguide in the Injector has been taken down, I will continue to disconnect cabling from module next week."

VBV2C20 Valve Interlock Bypass, ABIL tag #6209, Expected Removal Date changed.

Recent Log Entries -

Log Number: 4153394Logbook(s): Body: The following Bypassed Interlock had it's Expected Removal Date changed in the Accelerator Bypassed-Interlocks Log (ABIL) Previous Expected Removal Date: 31 Mar 2023 New Expected Removal Date: 31 May 2023 Who made the change: Mathew Almodovar-Mercado Justification for the change: Start date pushed back until other work during SAD is complete.<

PSS Firmware Updates continuing

Recent Log Entries -

Log Number: 4153368Logbook(s): Body: 

PSS firmware updates continuing in the MCC and progressing well.
Completed checks of Injector and NL segments.
Also updated the Firewall to permit OPC/UA traffic for future use.
On Friday will complete Hall A, Hall B and Hall C segments.

status phase2 injector work Day 4

Recent Log Entries -

Log Number: 4153337Logbook(s): Body: 

finished disconnecting all water lines from phase2 beamline, plugged the supply and return lines at the manifold, water back on, no leaks
Disconnected the water to buncher (ball valves inside the tunnel, along cable tray above chopper). Now there’s a leak at one of the supply lines at chopper2 cavity. Small leak, I turned OFF water flow.
All magnets and I&C stuff now disconnected
Larry removed most of the waveguide, thanks
Prepped the removal of the spectrometer beamline for tomorrow

Work preformed in the SAD

Recent Log Entries -

Log Number: 4153306Logbook(s): Entry Makers: lfarrishBody: 

This morning replaced all waveguide air supply hoses in the new zones NL21-26, except the main supply at the top of the ceiling I will get a taller ladder for that next week.
Waveguide in the Injector has been taken down, I will continue to disconnect cabling from module next week.

COMMENT ON Install and HV process 200 kV electron gun

Recent Log Entries -

Log Number: 4153301Logbook(s): Body: 

Particle counter shows 50-100 inside the clean hood.
Krypton evacuated between the bottle regulator and leak valve at the gun. I left the Kr line under vacuum for now.
After evacuating Kr line, connected pump cart to right angle valve upstream of gun overboard IP. Right angle valve remains closed of course. Started pumping on the line at ~ 11 am.
All of these tasks are in preparation for venting and removing HV chamber on Monday 03/27/23.

Pages

Subscribe to cebaf2013 aggregator